Two buses leave the same station at 7:00PM. One bus travels north at the rate of 30 kph and the other travels east at 40 kph. How many km apart are the buses at 8:00PM?
Answer by mananth(16949)

After 1 hour the buses are 30 km and 40 km away from the starting point.

they are travelling at right angles.
At any given time the starting point and the buses form a right angle
The distance between the buses will be the hypotenuse.
Using Pythagoras theorem
hypotenuse ^2 = side1^2 +side2^2


= 30^2 +40^2
= 900+1600
=2500
take the square root
hypotenuse = 25 km
They are 25 km away from each other